40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 26 17:05:44 2009, rob, Update, PSL, MOPA re-activated DSC_0513.JPGDSC_0517.JPG
    Reply  Tue May 26 17:34:14 2009, rob, Update, PSL, MOPA re-deactivated laser_temp.jpg
       Reply  Wed May 27 10:54:09 2009, rob, Update, PSL, we don't understand the chiller (broken) 
          Reply  Wed May 27 15:59:44 2009, rob, Update, PSL, we don't understand the chiller (broken) 
             Reply  Thu May 28 14:34:25 2009, rob, Update, PSL, chiller diagnosis 
Message ID: 1627     Entry time: Wed May 27 10:54:09 2009     In reply to: 1626     Reply to this: 1628
Author: rob 
Type: Update 
Category: PSL 
Subject: we don't understand the chiller (broken) 

Quote:

Quote:

steve, rob, alberto

 

Steve installed two rotary flow meters into the MOPA chiller system--one at the chiller flow output and one in the NPRO cooling line.  After some hijinks, we discovered that the long, insulated chiller lines have the same labels at each end.  This means that if you match up the labels at the chiller end, at the MOPA end you need switch labels: out goes to in and vice-versa.  This means that, indubitably, we have at some point had the flow going backwards through the MOPA, though I'm not sure if that would make much of a difference. 

Steve also installed a new needle valve in the NPRO cooling line, which works as expected as confirmed by the flow meter. 

We also re-discovered that the 40m procedures manual contains an error.  To turn on the chiller in the MOPA start-up process, you have to press ON, then RS-232, then ENTER.  The proc man says ON, RS-232, RUN/STOP.

The laser power is at 1.5W and climbing.

 Rob, Alberto

The chiller HT alarm started blinking, as the water temperature had reached 40 degrees C, and was still rising.  We turned off the MOPA and the chiller.  Maybe we need to open the needle valve a bit more?  Or maybe the flow needs to be reversed?  The labels on the MOPA are backwards?

 The chiller appears to be broken.  We currently have it on, with both the SENSOR and RS-232 unplugged.  It's running, circulating water, and the COOL led is illuminated.  But the temperature is not going down.  The exhaust out the back is not particularly warm.  We think this means the refrigeration unit has broken, or the chiller computer is not communicating with the refrigerator/heat exchanger.  Regardless, we may need a new chiller and a new laser.

ELOG V3.1.3-